List the elements of the given set in roster notation
{x|7-x=9;x,an integer} thank you

Answer by jim_thompson5910(35256)

7-x = 9

7 = 9 + x

7 - 9 = x

-2 = x

x = -2

So the set

{x|7-x=9;x,an integer}

is the same as the set

{ -2 }

Since only the number -2 satisfies the equation 7 - x = 9